OpenStudy (anonymous):

what is the answer of 6/2(1+2)

OpenStudy (anonymous):

9

OpenStudy (anonymous):

9 Step 1: 1+2 = 3 Step 2: 6/2 = 3 Step 3: 3*3 = 9

OpenStudy (anonymous):

use BODMAS - Bracket Of Division Multiplication Addition Subtraction in this order only try solving the question